Consider the probability distribution shown below.

x 0 1 2
P(x) 0.45 0.20 0.45

(a) Compute the expected value of the distribution.

(b) Compute the standard deviation of the distribution


Compute the standard deviation of the distribution. (Round your answer to four decimal places.)
Answer #1

Given data is false

because \Sigma p(x)=0.45+0.20+0.45=1.10\neq 1

if our assumption is true then

(a) E(x)=x.P(x)=0(0.45)+1(0.20)+2(0.45)=0+0.2+0.9=1.1

(b)\sigma ^{2}=\Sigma x^{2}.P(x)=0(0.45)+1(0.20)+4(0.45)-1.1^{2}=0.79

Standard deviation \sigma =\sqrt{0.79}=0.8888